OpenStudy (anonymous):

Is there a relatively compact way of writing the solution to: \[ \int_ \frac{\cos(x)^n}{\sin(x)^m} dx \] and \[ \int_ \frac{\cos(x)^n}{\sin(x)^m} dx \] ?

OpenStudy (experimentx):

what's that sin^n(x) or sin(x^n)?

OpenStudy (anonymous):

\[\int\limits\frac{(\cos(x))^n}{(\sin(x))^m}\]\[\int\limits\frac{(\cos(x))^m}{(\sin(x))^n}\]

OpenStudy (experimentx):

is it indefinite or are there limits?

OpenStudy (anonymous):

Indefinite

OpenStudy (experimentx):

Damn!! this isn't nice ... you assume m > n or n > m and use integration by parts n or m times ...

OpenStudy (experimentx):

also changing sine and cosine into complex form might work ... .

OpenStudy (anonymous):

That's weird, you'd have the imaginary part of e^ix above and real part below, or vice-versa.

OpenStudy (experimentx):

yeah ... just take Real part ... ignore imaginary part.

OpenStudy (anonymous):

Does that mean leave sin as sin?

OpenStudy (anonymous):

 If you're doing cos/sin